P6-25 (similar to) 20 2 F12 Del 0
Answer #1
When the interest is being paid out half yearly, then interest rates got half and periods double
VALUE OF BOND:
FORMULA = COUPON * PVIFA(YIELD REQD. , YEARS TO MATURITY) + PAR VALUE * PVIF(YIELD REQD. , YEARS TO MATURITY)
BOND A = (500* 9% / 2) * PVIFA(4% , 18) + 500 * PVIF(4% , 18) = 22.5 * 12.659 + 500 * 0.494 = $531.83
BOND B = (500* 11% / 2) * PVIFA(6.5% , 40) + 500 * PVIF(6.5% , 40) = 27.5 * 14.1455 + 500 * 0.0805 = $429.25
BOND C = (500* 14% / 2) * PVIFA(8% , 8) + 500 * PVIF(8% , 8) = 35 * 5.7466 + 500 * 0.5403 = $471.28